OpenStudy (anonymous):

find the circumference of a circle whose area is 60 pie cm squared. a. 2 square root of 15 pie cm b. 4 square root of 15 pie cm c. 30 pie cm d. 900 pie cm

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

PLZZ HELP

OpenStudy (anonymous):

@wolf1728

OpenStudy (anonymous):

circumference of a circle =2*pi*r area of a circle =pi*r^2 solve of r => r=sqrt(area/pi)

OpenStudy (anonymous):

i got 376

OpenStudy (anonymous):

60 pi=pi r^2 => r=sqrt(60)

OpenStudy (anonymous):

so whats the equation??? im lost

OpenStudy (anonymous):

first find r using the area equation =pi r^2then use the value of r in circumference equatio =2*pi*r

OpenStudy (anonymous):

whats r 60??

OpenStudy (wolf1728):

r = sq root of 60

OpenStudy (anonymous):

7.745?

OpenStudy (wolf1728):

Yes - actually it is 7.746

OpenStudy (anonymous):

7.746 times 2 and pi?? is that the final answer after??

OpenStudy (wolf1728):

you need the circumference circum = 2 * PI * radius

OpenStudy (anonymous):

48

OpenStudy (anonymous):

r=sqrt(60)=sqrt(4*16)=2sqrt(15) circ=2*pi*r= 2*2sqrt(15)*pi= 4 sqrt(15) pi

OpenStudy (anonymous):

so its b? the final answer???

OpenStudy (wolf1728):

I get 48.67

OpenStudy (anonymous):

is the final answer 4 square root of 15?

OpenStudy (anonymous):

yes

OpenStudy (wolf1728):

I get 7.746*2 PI or 15.49 PI

OpenStudy (anonymous):

ty

OpenStudy (wolf1728):

yes, that's it 4 * sq root (15)

OpenStudy (anonymous):

thx guys i dont know who to give the medal to